OpenStudy (anonymous):

How do you convert degrees to radians? example 145 degrees

OpenStudy (anonymous):

cool name man

OpenStudy (anonymous):

\[radians=\frac{ degrees*\pi }{ 180 }\]

OpenStudy (mertsj):

Why can't you use the same conversion factor I gave you for the last problem you posted?

OpenStudy (anonymous):

I'm new at this, chill

OpenStudy (anonymous):

just use the formula i posted up and you will be fine

OpenStudy (anonymous):

thank you galacticwaveXX

OpenStudy (mertsj):

You'll be fine except for the small problem that that formula is wrong.

OpenStudy (mertsj):

The conversion factor is : \[\frac{\pi radians}{180 degrees}\]

OpenStudy (anonymous):

Really can you explain again?

OpenStudy (anonymous):

Ok thanks, I'll try to figure that out and confirm

OpenStudy (mertsj):

If you are converting degrees to radians, multiply by the fraction I posted.

OpenStudy (mertsj):

If you are converting radians to degrees, multiply by the reciprocal of the fraction I posted.

OpenStudy (anonymous):

did you get the right answer/

OpenStudy (anonymous):

I know it equals 29pi over 36 but can you draw it out? I'm a visual learner and I just can't seem to wrap my head around this problem.

OpenStudy (anonymous):

you don't need to draw anything...all you need to do is plug it into the equation you should have done it like \[\frac{ 145 * \pi }{ 180 }=\frac{ 29 }{30 }\pi\] that's all you have to do

OpenStudy (anonymous):

that should be 29/36 my error but you get the idea

OpenStudy (anonymous):

My teacher wants us to show our work so I figured I could write it out

OpenStudy (mertsj):